# GaussianFlowOcc: Sparse and Weakly Supervised Occupancy Estimation using Gaussian Splatting and Temporal Flow

### v2

@@ 1 Introduction
Vision-based 3D object detection and occupancy estimation are fundamental challenges in the field of autonomous driving, as an accurate understanding of a vehicle’s surroundings is essential for reliable planning and navigation [60, 45, 70]. Despite recent advancements, existing approaches face several significant limitations. First, most recent occupancy estimation models rely on fully annotated 3D ground-truth data for training, which is expensive and challenging to obtain at scale. Second, these models depend on dense 3D voxel-based representations, leading to substantial computational overhead that is often unnecessary, given the natural sparsity of real-world 3D scenes. Although self-supervised methods have been developed to reduce reliance on 3D labels by leveraging temporal rendering [20, 67], these approaches fail to account for the inherent dynamics of autonomous driving scenarios. This introduces temporal inconsistencies during training, severely limiting their performance.
@@
To address these limitations, we introduce GaussianFlowOcc, a novel occupancy estimation method that combines efficiency and expressiveness through several contributions. At the core of our approach is a sparse representation of the scene that utilizes 3D Gaussian distributions, replacing traditional dense voxel grids. We propose a Gaussian Transformer network, a novel deep learning model designed to directly transform image features into 3D Gaussians. This shift not only reduces computational overhead, but also provides a more flexible representation for modeling dynamic scenes. Additionally, it allows us to leverage Gaussian Splatting [25] for efficient training without relying on 3D voxel labels. The estimated Gaussians are rendered back into the input camera space, where losses can be computed using 2D labels such as pseudo ground truth derived from pretrained models like Grounded-SAM [44] for semantics and Metric3D [63] for depth. Another key contribution is Temporal Gaussian Splatting, where adjacent frames are rendered to enhance the supervisory signal, allowing the model to capture 3D geometry more effectively. To handle temporal inconsistencies caused by dynamic objects, GaussianFlowOcc extends this framework with a Temporal Module that estimates a 3D flow for each Gaussian to each rendered temporal frame. This additional modeling of scene dynamics mitigates errors caused by object motion during training, strengthening the supervision signal and improving the overall performance. By combining these techniques, GaussianFlowOcc achieves state-of-the-art performance while maintaining computational efficiency, rendering it highly effective for practical occupancy estimation applications. The source will be made available at https://github.com/boschresearch/GaussianFlowOcc. In summary, our contributions are:
@@
  • Efficient Gaussian Transformer for sparse 3D Gaussian scene representation: We introduce a novel Gaussian Transformer that leverages set-based attention mechanisms to efficiently estimate a sparse set of 3D Gaussian distributions representing the scene from multi-view images. It entirely eliminates voxelization and computationally intensive operations, leading to an inference speedup of 50×50$\times$50 × compared to state-of-the-art.
  • Weakly supervised training with Gaussian Splatting: By leveraging Gaussian Splatting, our method can learn from 2D pseudo labels instead of relying on costly 3D annotations from other sensors like LiDAR. Our approach requires only images for training, which is the essential input needed for inference anyway, thus making it more scalable than methods depending on specialized 3D labeled datasets.
@@
  • Modeling of scene dynamics via learned 3D flow: GaussianFlowOcc integrates a Temporal Module that additionally estimates 3D flow for each Gaussian, enabling robust handling of dynamic objects during inference and during Temporal Gaussian Splatting at training time. GaussianFlowOcc is the first method to incorporate temporal supervision while explicitly modeling scene dynamics, without requiring ground-truth flow annotations.
@@ 2.1 Occupancy Estimation
The 3D semantic occupancy estimation task has become an important area in the autonomous driving research community in recent years. Numerous voxel-based occupancy benchmarks have been introduced for datasets such as SemanticKITTI [3] and nuScenes [52, 57]. Early works in 3D occupancy estimation utilize established techniques from Birds-Eye-View (BEV) perception and object detection [19, 29] to lift multi-view camera images into a unified 3D voxel grid [21, 18, 53, 7, 28]. Subsequent approaches have improved model efficiency [65, 55, 34, 31, 46, 51, 22], optimized the training procedure and label efficiency [40, 5, 14, 16, 48], and improved occupancy estimation performance through architectural innovations [30, 69, 23, 49, 71, 36, 35].
Despite these advancements, the majority of existing 3D occupancy estimation methods require costly voxel-based 3D ground truth labels for training. Therefore, a parallel line of research has emerged that focuses on self- and weakly supervised learning for occupancy estimation models, leveraging only 2D labels. SelfOcc [20] and OccNeRF [67] employ volume rendering inspired by NeRF [37] to render estimated 3D occupancy back to the 2D image space. Spatial and temporal photometric consistency losses can be used to train the geometry estimation, while semantic information is incorporated using pretrained vision foundation models like OpenSeeD [68] or GroundedSAM [44]. GaussianOcc [15] replaces the volume rendering pipeline with 3D Gaussian Splatting to accelerate training, yet it continues to model the scene with dense occupancy grids, thus failing to exploit the benefits of a fully sparse Gaussian representation. Also, while these methods eliminate the need for 3D labels, they fail to address scene dynamics, a critical aspect when relying on temporal consistency losses.
@@ 2.2 Differentiable Rendering and 3D Gaussian Splatting
Differentiable rendering has emerged as a powerful technique for learning 3D scene representations by projecting them into 2D views, followed by an optimization based on photometric or semantic consistency. Neural Radiance Fields (NeRF) [37] have been particularly influential, modeling scenes as volumetric representations that encode radiance and density, enabling novel view synthesis through differentiable volume rendering. Recently, 3D Gaussian Splatting (GS) [25] has introduced a novel paradigm for 3D scene reconstruction by representing scenes as a collection of 3D Gaussians. This approach significantly reduces computational overhead while preserving expressive scene modeling, making it highly efficient for tasks requiring dynamic or large-scale reconstructions. NeRF and GS approaches were originally designed to reconstruct individual scenes for novel-view synthesis, with research focusing on improving efficiency [38, 11], rendering quality [1, 2] or feature enrichment [42, 26, 62, 75]. Several works have further explored the modeling of dynamic scenes for video reconstruction [59, 13, 41]. Another line of work incorporates different priors like depth, stereo-matching, or additional data like LiDAR, to train generalizable reconstruction models [61, 9, 10, 64, 58, 32, 73]. As mentioned above, similar ideas have been adapted to train occupancy estimation models. Methods like OccNeRF [67], SelfOcc [20] and GaussianOcc [15] employ volume rendering or GS for weakly supervised training. However, they still rely on voxel grids for scene representation, which limits efficiency and scalability. GSRender [48] additionally introduces a ray compensation method to mitigate duplicate predictions along camera rays, a common issue among rendering-based approaches. Lastly, the recent approach GaussTR [24] shares similarities with our method by adopting 3D Gaussians as the scene representation. However, GaussTR uses multiple pretrained feature encoders (e.g., CLIP, Metric3D, SAM) during inference, making the pipeline computationally expensive. Furthermore, GaussTR employs standard attention layers, which significantly limits the number of Gaussians it can handle.
@@ 3.1 Problem Definition and Scene Representation
The objective of occupancy prediction is to accurately estimate the 3D geometry and semantics surrounding a vehicle based on a set of $L$ multi-view images $I=\{I^{1},I^{2},...,I^{L}\}$ at the current time step. Previous methods have typically utilized a semantic voxel volume $V=\{c_{1},c_{2},...c_{C}\}^{X\times Y\times Z}$ on a predefined grid, with $C$ representing the number of semantic classes, as a structured representation of the scene. In contrast, our approach involves estimating the 3D scene using 3D Gaussian distributions, drawing inspiration from 3D Gaussian Splatting [25]. The scene is defined as a collection of Gaussians $\mathcal{G}=\{G_{1},G_{2},...,G_{n}\}$, each characterized by a set of properties: mean $\mu\in\mathbb{R}^{3}$, opacity $\sigma\in[0,1]$, scale $s\in\mathbb{R}^{3}$, rotation quaternions $r\in\mathbb{R}^{4}$, and semantic logits $c\in\mathbb{R}^{C}$:
$G_{i}=(\mu,\sigma,s,r,c)$ (1)
@@ 3.2.2 Image Encoder
Image features $\hat{I}$ are extracted from the input images $I$ using a pretrained backbone architecture like ResNet50 [17]. These features and the initial Gaussians are then fed into the Gaussian Transformer.
@@ 3.2.3 Gaussian Transformer
Figure 3: Architecture of the proposed Gaussian Transformer. In each block, the features and means of the Gaussians are refined using a Positional Encoding, ITA, ISA, GICA and Gaussian Rectification.
@@
To enable Gaussians to acquire meaningful scene information, we introduce the Gaussian-Image Cross-Attention (GICA), which facilitates interactions between Gaussians $\mathcal{G}$ and the image features $\hat{I}$. For this we leverage deformable cross-attention, originally proposed by [29] and widely adopted in subsequent works. In this operation, the current positions of the Gaussians $\mathcal{G}_{\mu}$ are projected onto the image feature maps using the camera parameters. For each projected point, a set of image features around this point are sampled and used as keys and values for an attention operation. This allows the Gaussians to extract rich, localized information from the scene in a computationally efficient manner. For a detailed description of deformable attention, we refer readers to [29].
@@
Computing interactions among Gaussians is a critical aspect of our approach. However, the standard attention mechanism is characterized by quadratic time and memory complexity $\mathcal{O}(N^{2})$. This imposes severe constraints on the number of Gaussians that can be processed (as demonstrated in Fig. 5). To address this limitation, we draw inspiration from the Set Transformer [27] and employ an attention operation called Induced Self-Attention (ISA). ISA reduces the memory complexity to approximately linearly scaling with respect to the number of Gaussians, allowing us to employ a significantly larger number of Gaussians. The operation is schematically illustrated in Fig. 4 (A). In essence, ISA replaces quadratic attention with two attention operations connected via a bottleneck. Instead of allowing Gaussians to directly attend to one another, we introduce a set of $M$ trainable latent feature vectors, called inducing points $P\in\mathbb{R}^{M\times D}$ (with $M\ll N$). These inducing points aggregate information from all Gaussians into a bottleneck representation $H$. Subsequently, all Gaussians interact with the bottleneck features $H$ enabling indirect yet complete Gaussian-to-Gaussian interactions. This transformation reduces the computational complexity to $\mathcal{O}(MN)$. The overall process is formally defined as:
@@
Building upon the concept of ISA, we introduce Induced Temporal Attention (ITA) to enable efficient temporal information propagation across frames, as depicted in Fig. 4 (B). The ITA operation leverages the structure of induced attention by letting the inducing points first attend to the Gaussian features from the previous frame to compute the bottleneck features $Z$. Subsequently, Gaussian features from the current frame interact with the bottleneck features, facilitating temporal attention with linear complexity. Formally, given $\mathcal{G}^{\prime}_{f}$ as the final Gaussian features from the previous frame, the ITA operation is expressed as:
@@
$\ \text{where}\quad Z=\mathrm{MHA}(P,\mathcal{G}^{\prime}_{f},\mathcal{G}^{\prime}_{f}),$ (5)
Note that, as we show in Fig. 2 and Fig. 3, the previous Gaussian features $\mathcal{G}^{\prime}_{f}$ are derived by first correcting the previous final Gaussian means using ego-motion and estimated temporal flow to the current frame, computing the positional encoding and adding them to the previous final features.
@@
$\ \mathcal{G}^{b+1}_{\mu}=\mathcal{G}^{b}_{\mu}+\Delta\mathcal{G}^{b}_{\mu}\;$ (6)
$\ \text{where}\quad\Delta\mathcal{G}^{b}_{\mu}=\mathrm{Rect}(\mathcal{G}^{b}_{f})$ (7)
@@ 3.2.4 Gaussian Heads
$\ r=\mathrm{norm}(\mathrm{MLP}_{r}(\mathcal{G}_{f})),\ c=\mathrm{MLP}_{c}(\mathcal{G}_{f}),$ (9)
@@ 3.3 Gaussian Splatting Supervision
Representing the scene with 3D Gaussians allows us to leverage the highly efficient Gaussian Splatting (GS) [25] method during training. For a detailed explanation of GS, we refer readers to the original work [25]. Essentially, the estimated Gaussian properties are used to project the 3D Gaussians back onto the input image views via the corresponding camera parameters. This enables us to rasterize predicted 2D semantic logits and depth values. Notably, unlike standard GS, which rasterizes RGB values, we rasterize semantic logits instead. We generate 2D semantic and depth labels using pretrained models to serve as ground truth. The entire model is then trained using 2D rendering losses. This enables training in a weakly supervised manner directly in the 2D image space, eliminating the need for 3D labels or additional sensor data, such as LiDAR, entirely. We optimize the MSE-Loss $\mathcal{L}_{depth}$ between rendered depth $\hat{D}$ and precomputed depth maps $D$ and the binary cross-entropy loss $\mathcal{L}_{seg}$ between rendered logits $\hat{S}$ and precomputed semantic segmentation maps $S$.
$\mathcal{L}=\mathcal{L}_{depth}(\hat{D},D)+\mathcal{L}_{seg}(\hat{S},S)$ (10)

### v3

@@ 1 Introduction
Vision-based 3D object detection and occupancy estimation are fundamental challenges in the field of autonomous driving, as an accurate understanding of a vehicle’s surroundings is essential for reliable planning and navigation [62, 46, 73]. Despite recent advances, existing approaches face several limitations. First, most recent occupancy estimation models rely on fully annotated 3D ground truth data for training, which is expensive and challenging to obtain at scale. Second, these models depend on dense 3D voxel-based representations, leading to substantial computational overhead that is often unnecessary, given the natural sparsity of real-world 3D scenes. Although self-supervised methods have been developed to reduce reliance on 3D labels by leveraging temporal rendering [20, 70], these approaches fail to account for the inherent dynamics of driving scenarios. This introduces temporal inconsistencies during training, severely limiting their performance.
@@
To address these limitations, we introduce GaussianFlowOcc, a novel occupancy estimation method that combines efficiency and expressiveness through several contributions. At the core of our approach is a sparse representation of the scene that utilizes 3D Gaussian distributions, replacing traditional dense voxel grids. We propose a Gaussian Transformer network, a novel deep learning model designed to directly transform image features into 3D Gaussians. This shift not only reduces computational overhead but also provides a more flexible representation for modeling dynamic scenes. Additionally, it allows us to leverage Gaussian Splatting [25] for efficient training without relying on 3D voxel labels. The estimated Gaussians are rendered back into the input camera space, where losses can be computed using 2D labels such as pseudo ground truth derived from pretrained models like Grounded-SAM [45] for semantics and Metric3D [66] for depth. Another key contribution is Temporal Gaussian Splatting, where adjacent frames are rendered to enhance the supervisory signal, allowing the model to capture 3D geometry more effectively. To handle temporal inconsistencies caused by dynamic objects, GaussianFlowOcc extends this framework with a Temporal Module that estimates a 3D flow for each Gaussian to each rendered temporal frame. This additional modeling of scene dynamics mitigates errors caused by object motion during training, strengthening the supervision signal and improving the overall performance. By combining these techniques, GaussianFlowOcc achieves state-of-the-art performance while maintaining computational efficiency, rendering it highly effective for practical occupancy estimation applications. In summary, our contributions are:
@@
  • Efficient Gaussian Transformer for sparse 3D Gaussian scene representation: We introduce a novel Gaussian Transformer that leverages set-based attention mechanisms to efficiently estimate a sparse set of 3D Gaussian distributions representing the scene from multi-view images. It entirely eliminates voxelization and computationally intensive operations, leading to an inference speedup of $50\times$ compared to state-of-the-art. At the same time, by leveraging Gaussian Splatting, our method learns from 2D pseudo labels instead of relying on costly 3D annotations from other sensors like LiDAR. Our approach requires only images for training, making it more scalable than methods depending on 3D labeled datasets.
@@
  • Modeling of scene dynamics via learned 3D flow: GaussianFlowOcc integrates a Temporal Module that additionally estimates 3D flow for each Gaussian, enabling robust handling of dynamic objects during inference and during Temporal Gaussian Splatting at training time. GaussianFlowOcc is the first method to incorporate temporal supervision while explicitly modeling scene dynamics, without requiring ground truth flow annotations.
@@ 2.1 Occupancy Estimation
The 3D semantic occupancy estimation task has become an important area in the autonomous driving research community in recent years. Numerous voxel-based occupancy benchmarks have been introduced for datasets such as SemanticKITTI [3] and nuScenes [53, 58]. Early works in 3D occupancy estimation utilize established techniques from Birds-Eye-View (BEV) perception and object detection [19, 29] to lift multi-view camera images into a unified 3D voxel grid [21, 18, 54, 7, 28]. Subsequent approaches have improved model efficiency [68, 56, 35, 31, 47, 52], optimized the training procedure and label efficiency [41, 5, 14, 16, 49], and improved occupancy estimation performance through architectural innovations [30, 72, 23, 50, 74, 37, 36]. Specifically, GaussianFormer [22] and GaussianWorld [81] employ a sparse and efficient Gaussian representation similar to our work.
Despite these advancements, the majority of existing 3D occupancy estimation methods require costly voxel-based 3D ground truth labels for training. Therefore, a parallel line of research has emerged that focuses on self and weakly supervised learning for occupancy estimation models, leveraging only 2D labels. SelfOcc [20] and OccNeRF [70] employ volume rendering inspired by NeRF [38] to render estimated 3D occupancy back to the 2D image space. Photometric consistency losses can be used to train the geometry estimation, while semantic information is incorporated using pretrained foundation models like OpenSeeD [71] or GroundedSAM [45]. GaussianOcc [15] replaces the volume rendering pipeline with 3D Gaussian Splatting to accelerate training, yet it continues to model the scene with dense occupancy grids, thus failing to exploit the benefits of a fully sparse Gaussian representation. Also, while these methods eliminate the need for 3D labels, they fail to address scene dynamics, a critical aspect when relying on temporal consistency losses.
@@ 2.2 Differentiable Rendering and 3D Gaussian Splatting
Differentiable rendering has emerged as a powerful technique for learning 3D scene representations by projecting them into 2D views, followed by an optimization based on photometric or semantic consistency. Neural Radiance Fields (NeRF) [38] have been particularly influential, modeling scenes as volumetric representations that encode radiance and density, enabling novel view synthesis through differentiable volume rendering. Recently, 3D Gaussian Splatting (GS) [25] has introduced a novel paradigm for 3D scene reconstruction by representing scenes as a collection of 3D Gaussians. This approach significantly reduces computational overhead while preserving expressive scene modeling, making it highly efficient for dynamic or large-scale reconstructions. NeRF and GS approaches were originally designed to reconstruct individual scenes for novel-view synthesis, with research focusing on improving efficiency [39, 11], rendering quality [1, 2] or feature enrichment [43, 26, 65, 79]. Several works have further explored the modeling of dynamic scenes for video reconstruction [61, 13, 42]. Another line of work incorporates different priors like depth, stereo-matching, or additional data like LiDAR, to train generalizable reconstruction models [63, 9, 10, 67, 60, 32, 76]. Several methods have also been developed specifically for reconstruction within autonomous driving scenarios [34, 80, 64], while GaussianAD [77] employs a 3D Gaussian representation for end-to-end driving. As mentioned above, similar ideas have been adapted to train occupancy estimation models. Methods like OccNeRF [70], SelfOcc [20] and GaussianOcc [15] employ volume rendering or GS for weakly supervised training. However, they still rely on voxel grids for scene representation, which limits efficiency and scalability. GSRender [49] additionally introduces a ray compensation method to mitigate duplicate predictions along camera rays. Lastly, the recent approach GaussTR [24] shares similarities with our method by adopting 3D Gaussians as the scene representation. However, GaussTR uses multiple pretrained feature encoders (e.g., CLIP, Metric3D, SAM) during inference, making the pipeline computationally expensive. Furthermore, GaussTR employs standard attention layers, which significantly limits the number of Gaussians it can handle.
@@ 3.1 Problem Definition and Scene Representation
The objective of occupancy prediction is to accurately estimate the 3D geometry and semantics surrounding a vehicle based on a set of $L$ multi-view images $I=\{I^{1},I^{2},...,I^{L}\}$ at the current time step. Previous methods have typically utilized a semantic voxel volume $V=\{c_{1},c_{2},...c_{C}\}^{X\times Y\times Z}$ on a predefined grid, with $C$ representing the number of semantic classes, as a structured representation of the scene. In contrast, our approach involves estimating the 3D scene using 3D Gaussian distributions, drawing inspiration from 3D Gaussian Splatting [25]. The scene is defined as a collection of $N$ Gaussians $\mathcal{G}=\{G_{1},G_{2},...,G_{N}\}$, each characterized by a set of properties: mean $\mu\in\mathbb{R}^{3}$, opacity $\sigma\in[0,1]$, scale $s\in\mathbb{R}^{3}$, rotation quaternions $r\in\mathbb{R}^{4}$, and semantic logits $c\in\mathbb{R}^{C}$:
$G_{i}=(\mu,\sigma,s,r,c).$ (1)
@@ 3.2.2 Image Encoder
Image features $\hat{I}$ are extracted from the input images $I$ using a pretrained backbone architecture like ResNet-50 [17]. These features and the initial Gaussians are then fed into the Gaussian Transformer.
@@ 3.2.3 Gaussian Transformer
Figure 3: Architecture of the proposed Gaussian Transformer. In each block, the features and means of the Gaussians are refined using Pos Enc, ITA, ISA, GICA and Gaussian Rectification.
@@
To enable Gaussians to acquire meaningful scene information, we introduce the Gaussian-Image Cross-Attention (GICA), which facilitates interactions between Gaussians $\mathcal{G}$ and the image features $\hat{I}$. For this we leverage deformable cross-attention, originally proposed by [29] and widely adopted in subsequent works. In this operation, the current positions of the Gaussians $\mathcal{G}_{\mu}$ are projected onto the image feature maps using the camera parameters. For each projected point, a set of image features around this point is sampled and used as keys and values for an attention operation. This allows the Gaussians to extract rich, localized information from the scene in a computationally efficient manner. For a detailed description of deformable attention, we refer readers to [29].
@@
Computing interactions among Gaussians is a critical aspect of our approach. However, the standard attention mechanism is characterized by quadratic time and memory complexity $\mathcal{O}(N^{2})$. This imposes severe constraints on the number of Gaussians that can be processed (as demonstrated in Fig. 5). To address this limitation, we draw inspiration from the Set Transformer [27] and employ an attention operation called Induced Self-Attention (ISA). ISA reduces the memory complexity to approximately linearly scaling with respect to the number of Gaussians, allowing to employ a significantly larger number of Gaussians. The operation is schematically illustrated in Fig. 4 (A). In essence, ISA replaces quadratic attention with two attention operations connected via a bottleneck. Instead of allowing Gaussians to directly attend to one another, we introduce a set of $M$ trainable latent feature vectors, called inducing points $P\in\mathbb{R}^{M\times D}$ (with $M\ll N$). These inducing points aggregate information from all Gaussians into a bottleneck representation $H$. Subsequently, all Gaussians interact with the bottleneck features $H$ enabling indirect yet complete Gaussian-to-Gaussian interactions. This transformation reduces the computational complexity to $\mathcal{O}(MN)$. The overall process is formally defined as:
@@
Building upon the concept of ISA, we introduce Induced Temporal Attention (ITA) to enable efficient temporal information propagation across frames, as depicted in Fig. 4 (B). The ITA operation leverages the structure of induced attention by letting the inducing points $P$ first attend to the Gaussian features from the previous frame to compute the bottleneck features $Z$. Subsequently, Gaussian features from the current frame interact with the bottleneck features, facilitating temporal attention with linear complexity. Formally, given $\mathcal{G}^{\prime}_{f}$ as the final Gaussian features from the previous frame, the ITA operation is expressed as:
@@
$\ \text{where}\quad Z=\mathrm{MHA}(P,\mathcal{G}^{\prime}_{f},\mathcal{G}^{\prime}_{f}).$ (5)
@@
$\ \mathcal{G}^{b+1}_{\mu}=\mathcal{G}^{b}_{\mu}+\Delta\mathcal{G}^{b}_{\mu},\;$ (6)
$\ \text{where}\quad\Delta\mathcal{G}^{b}_{\mu}=\mathrm{Rect}(\mathcal{G}^{b}_{f}).$ (7)
@@ 3.2.4 Gaussian Heads
$\ r=\mathrm{norm}(\mathrm{MLP}_{r}(\mathcal{G}_{f})),\ c=\mathrm{MLP}_{c}(\mathcal{G}_{f}).$ (9)
@@
3.3 Temporal Propagation
As we show in Fig. 2 and Fig. 3, the previous Gaussian features $\mathcal{G}^{\prime}_{f}$ are derived by first correcting the previous final Gaussian means using ego-motion and estimated temporal flow to the current frame, computing the positional encoding and adding them to the previous final features. The ITA module then fuses the temporal information in a recurrent manner. The model always operates on inputs from a single time step while maintaining a memory of the last prediction to facilitate effective temporal fusion.
@@ 3.4 Gaussian Splatting Supervision
Representing the scene with 3D Gaussians allows us to leverage the highly efficient Gaussian Splatting (GS) [25] method during training. For a detailed explanation of GS, we refer to the original work [25]. Essentially, the estimated Gaussian properties are used to project the 3D Gaussians back onto the input image views via the corresponding camera parameters. This enables us to rasterize predicted 2D semantic logits and depth values. Notably, unlike standard GS, which rasterizes RGB values, we rasterize semantic logits instead. We generate 2D semantic and depth labels using pretrained models to serve as ground truth. The entire model is then trained using 2D rendering losses. This enables training in a weakly supervised manner directly in the 2D image space, eliminating the need for 3D labels or additional sensor data, such as LiDAR, entirely. We optimize the MSE-Loss $\mathcal{L}_{depth}$ between rendered depth $\hat{D}$ and precomputed depth maps $D$ and the binary cross-entropy loss $\mathcal{L}_{seg}$ between rendered logits $\hat{S}$ and precomputed semantic segmentation maps $S$:
$\mathcal{L}=\mathcal{L}_{depth}(\hat{D},D)+\mathcal{L}_{seg}(\hat{S},S).$ (10)